OpenStudy (anonymous):

Find all solutions in the interval [0,2pi) cos(x+pi/4)-cos(x-pi/4)=1

OpenStudy (anonymous):

\[\cos(x+\frac{ \Pi }{ 4})-\cos(x-\frac{ \Pi }{ 4})=1\]

OpenStudy (zehanz):

You could use the formula for the difference of two cosines:\[\cos a - \cos b = -2\sin \frac{a+b}{2}\sin \frac{a-b}{2}\]